# ViSTA-SLAM: Visual SLAM with Symmetric Two-view Association

### v1

@@ Abstract
{strip}
Figure 1: ViSTA-SLAM Results on a Multi-room Scene [8]. By combining the proposed lightweight frontend Symmetric Two-view Association (STA) model with $\mathrm{Sim(3)}$ pose graph optimization and loop closuring as the backend, ViSTA-SLAM achieves high-quality reconstruction and accurate trajectory estimation on challenging scenes while running in real time.
@@ 1.1 Real-time Monocular Dense SLAM
Simultaneous Localization and Mapping (SLAM) jointly estimates an agent’s pose and the surrounding 3D scene from sensor observations. In the monocular dense setting, a single RGB camera is used to reconstruct a continuous 3D map, enabling both geometric accuracy and visual realism. The camera poses and dense reconstruction outputs underpin downstream tasks [42, 2, 32, 31, 19] such as semantic perception, object interaction, and scene editing, and are crucial for applications in VR/AR, robotics, and autonomous driving, where accurate, low-latency 3D perception is essential.
@@ 1.2 Related Work
Classical Visual SLAM    Classical visual SLAM methods can be broadly categorized into two types. The first, similar to incremental SfM [1, 41, 26], is feature-based SLAM [10, 35, 34, 6], relying on keypoint extraction and descriptor matching to provide constraints for triangulation and PnP [22, 21] pose estimation. The second, known as direct methods, such as LSD-SLAM [15] and DSO [16], optimizes camera poses by minimizing photometric error from pixel intensities while estimating a per-frame depth map. Both categories typically adopt a frontend (feature-based or direct) and a backend for optimization, most often bundle adjustment [49] to jointly refine poses and structure. However, they rely heavily on accurate camera calibration and are generally limited to sparse 3D reconstructions.
@@
Dense Visual SLAM    To enable denser reconstructions, recent works have incorporated deep learning into either the frontend or the scene representation. For example, DROID-SLAM [48] uses RAFT [47] for dense optical flow in the frontend and performs dense bundle adjustment on the GPU, while methods such as SuperPrimitive [30] and COMO [12] leverage monocular priors (e.g., surface normals, depth distributions) from pretrained models with direct photometric optimization. BA-Track [7] augments point-based tracking with a scale-grid deformation of monocular depth priors. Neural scene representations have also been adopted, with NICER-SLAM [61] and MonoGS [29] optimizing camera poses and 3D structure using NeRF [33] and 3D Gaussian Splatting [20], respectively. Tracking robustness and geometric accuracy have been further improved by integrating depth priors and auxiliary trackers [57, 40, 60, 58, 59]. However, most approaches still require accurate camera intrinsics and many struggle to achieve true real-time performance due to the computational demands of dense optimization and neural rendering.
@@
SLAM with 3D Foundation Models    All aforementioned methods require known and accurate camera intrinsics. With the advent of 3D foundation models [53, 23, 51], several intrinsic-free SLAM frameworks have emerged, aiming to produce dense outputs without calibration. Methods such as Spann3R [50] and others [27, 52, 5] extend the two-view DUSt3R [53] model to sequential inputs, directly regressing point clouds in a unified global coordinate system. Reloc3r [13] instead regresses only relative poses and performs offline global optimization using SfM techniques [37, 46, 56]. MASt3R-SLAM [36] extracts dense correspondences from MASt3R [23] and feeds them into a classical optimization pipeline, while submap-based methods [28, 11] employs the multiview model VGGT [51] to regress local submaps before stitching them via pose graph optimization. While these approaches address some classical limitations, they still face notable drawbacks:
@@
    Current two-view models [53, 23] use asymmetric architectures that regress pointmaps of both views to the first view’s coordinates, making it difficult to decouple views for backend optimization (e.g. loop closure).
@@
    Pure regression methods [50, 52, 27] predict incoming frames with previous memory, but suffer from drift and start forgetting once the trajectory gets longer.
@@
    Methods [50, 52, 27, 36] built on current two-view models inherit the asymmetric architecture with two separate decoders, resulting in large model size. Submap-based methods [28, 11] employ an even larger multiview model [51] to build submaps, which further increases the size of the frontend model.
@@ 1.3 Contributions of ViSTA-SLAM
To address these concerns, we propose ViSTA-SLAM, a novel real-time monocular visual SLAM pipeline based on symmetric two-view association. At its core is a lightweight Symmetric Two-view Association (STA) model frontend, which takes two RGB images as input and simultaneously regresses two pointmaps in their respective local coordinate frames, along with the relative camera pose between them. During training, we enforce cycle consistency on relative poses and geometric consistency on pointmaps to improve accuracy and stability. Unlike prior 3D models [53, 23, 51], STA is fully symmetric with respect to its inputs: neither view is designated as a reference, and the same encoder–decoder architecture is applied to both. In the backend, we perform $\mathrm{Sim}(3)$ pose graph optimization with loop closures to mitigate drift and ensure global consistency. To further enhance robustness, each view is represented by multiple nodes rather than a single one, which are connected by scale-only edges to handle scale inconsistencies across different forward passes.
This symmetric design makes our frontend substantially more lightweight than existing methods, with STA being only 64% the size of MASt3R [23] and 35% the size of VGGT [51]. Unlike prior approaches [28, 11] that group multiple views into a single submap node, our method assigns each view its own nodes in the pose graph. Leveraging the local pointmaps produced by the STA frontend, each node can be represented independently, connected to others solely through relative transformations. Compared to submap-based methods [28], this design yields a more flexible graph structure and greater robustness. The combination of this flexibility and lightweight architecture underpins our choice of a symmetric two-view model as the frontend.
@@
    Our method achieves state-of-the-art performance on the real-world 7-Scenes [43] and TUM-RGBD [45] datasets for both camera trajectory estimation and dense 3D reconstruction.
@@ 2.1 Symmetric Two-view Association Model
Encoder    Our STA model takes in two images $\bm{I}_{i}$, $\bm{I}_{j}$ as input. It uses a shared ViT encoder [14] to divide input images into patches and encode them into features
@@
Symmetric Formulation    Prior approaches [53, 23] regress both point maps into the coordinate frame of the first view, thus requiring two separate decoders. In contrast, our model predicts only local point maps and relative poses between views. This fully symmetric formulation makes it possible to use only one decoder. As a result, the number of parameters for decoding is effectively reduced by half (shown in Fig. 4), forming a more compact model for real-time applications. Moreover, producing local view outputs in their own coordinate systems is better suited for the subsequent pose graph optimization, see Sec. 2.3 for details.
Local Point Maps    Given decoding features $\bm{D}_{i/j}^{(b)}$, we use a DPT head [38] to regress the local point maps $\bm{P}$ and corresponding confidence maps $\bm{W}$:
@@
Since $\bm{M}_{ij}$ is not guaranteed to lie on the $SO(3)$ manifold, we apply SVD orthogonalization [24] to it to obtain a valid rotation matrix $\bm{R}_{ij}$. Then, the relative transformation is $\bm{T}_{ij}=\left[\bm{R}_{ij}|\bm{t}_{ij}\right]$. With our symmetric formulation, we could also input $\bm{p}_{j}$ into the pose head to regress $\bm{T}_{ji}$. But in practice we only regress one of them for building pose graph.
@@ 2.2 Training Objective
Local Pointmap Loss    Following DUSt3R [53], we apply the confidence-weighted regression loss for all predicted point maps with valid ground truths. Since the reconstruction is up-to-scale, we also normalize the regressed pointmap and the ground-truth pointmap according to their mean Euclidean distance to the origin, $n$ and $\hat{n}$:
@@
Figure 3: Asymmetric vs. Symmetric Architectures. Asymmetric architectures [53, 23] use two decoders to regress point maps in a shared coordinate space. our symmetric formulation regresses relative pose and local point maps with only a single decoder, reducing over 36% of the parameters ($\sim$ 0.4 vs. 0.7 billion), while achieving higher accuracy and enabling pose graph optimization in the backend.
@@
NICER-SLAM [61] 0.033 0.069 0.042 0.108 0.200 0.039 0.108 0.086
DROID-SLAM [48] 0.036 0.027 0.025 0.066 0.127 0.040 0.026 0.049
Calib. GlORIE-SLAM [57] 0.036 0.029 0.014 0.094 0.144 0.053 0.020 0.056
CUT3R [52] 0.743 0.226 0.363 0.664 0.546 0.381 0.413 0.477
SLAM3R [27] 0.089 0.048 0.036 0.088 0.196 0.102 0.126 0.098
MASt3R-SLAM [36] 0.063 0.046 0.029 0.103 0.112 0.074 0.032 0.066
VGGT-SLAM [28] 0.037 0.026 0.022 0.103 0.147 0.063 0.095 0.070
@@
Table 1: Camera Trajectory Estimation (ATE RMSE) on 7-Scenes [43]. ViSTA-SLAM performs the best on average.
@@ 2.3 Backend Pose Graph Optimization
Loop Closure    We use Bag of Words [17] to detect loop candidates, which form new pairs. Then we can feed each candidates pair into our STA model to confirm these proximity. If the predicted confidence score of the relative pose is higher than a predefined threshold $\tau_{p}$, this pair is accepted as a valid loop, and two new nodes connected by a pose edge are added. Each new node is also connected to the first processed node of its corresponding view via a scale edge.
@@
ORB-SLAM3 [6] $\times$ 0.017 0.210 $\times$ 0.034 $\times$ $\times$ $\times$ 0.009 N/A
DPV-SLAM [25] 0.112 0.018 0.029 0.057 0.021 0.330 0.030 0.084 0.010 0.076
DPV-SLAM++ [25] 0.132 0.018 0.029 0.050 0.022 0.096 0.032 0.098 0.010 0.054
DROID-SLAM [48] 0.111 0.018 0.042 0.021 0.016 0.049 0.026 0.048 0.012 0.038
Calibrated GlORIE-SLAM [57] 0.128 0.016 0.028 0.021 0.021 0.042 0.020 0.035 0.010 0.036
CUT3R [52] 0.174 0.592 0.546 0.662 0.467 0.911 0.051 0.845 0.129 0.486
SLAM3R [27] 0.211 0.861 0.967 0.790 0.755 1.013 0.063 0.986 0.185 0.648
MASt3R-SLAM [36] 0.070 0.032 0.055 0.056 0.035 0.118 0.041 0.116 0.020 0.060
VGGT-SLAM [28] 0.063 0.031 0.048 0.152 0.023 0.133 0.038 0.039 0.020 0.061
@@
Table 2: Camera Trajectory Estimation (ATE RMSE) on TUM-RGBD [45]. ViSTA-SLAM performs the best on average.
@@ 3 Experiments
Evaluation Datasets and Metrics    Following VGGT-SLAM [28], we evaluate our method on standard monocular SLAM benchmarks for camera tracking accuracy and reconstruction quality. We report root mean square error (RMSE) of absolute trajectory error (ATE, in meters) on real-world 7-Scenes [43] and TUM-RGBD [45] datasets using the evo toolkit [18]. Reconstruction quality on 7-Scenes is assessed via RMSE of accuracy, completion, and Chamfer distance (meters), leveraging its ground-truth 3D scenes.
Implementation Details    The frontend STA model is initialized from the weights of DUSt3R [53], and trained on ScanNet [8], ScanNet++[54], ARKitScenes[4], CO3D [39], Aria Synthetic Environments [3], and Replica [44] for 7 days using 8 NVIDIA H100 GPUs. We use AdamW optimizer to train our STA model with learning rate $1.5e^{-5}$, weight decay $0.01$, $B=12$, $\alpha^{\text{point}}=0.2$, $\alpha^{\text{pose}}=0.05$, $\lambda_{\text{pmap}}=1$, $\lambda_{\text{pose}}=1$, $\lambda_{\text{gc}}=1$, and $\tau_{p}=0.75$. We conducted evaluations on a machine with an NVIDIA RTX 4090 GPU and an Intel i9-14900KF CPU, with $N=2$ for 7-Scenes and $N=3$ for TUM-RGBD.
Baselines    ViSTA-SLAM is primarily compared with state-of-the-art (SOTA) learning-based SLAM methods in uncalibrated scenarios: VGGT-SLAM [28], MASt3R-SLAM [36], SLAM3R [27], and CUT3R [52]. To reduce randomness from VGGT-SLAM’s RANSAC, we run it 5 times per scene as suggested in their paper; results for other methods, including ours, are deterministic. We also compare with SOTA methods using known camera intrinsics [6, 61, 57, 48, 25]. Some results are taken from [36, 28]. For MASt3R-SLAM and VGGT-SLAM, we keep their original keyframe selection; for ours, CUT3R, and SLAM3R, we use frame strides of 5 (7-Scenes) and 3 (TUM-RGBD). Calibrated methods are shown in gray. Best results are highlighted as first, second, and third.

### v2

@@ 1.1 Real-time Monocular Dense SLAM
Simultaneous Localization and Mapping (SLAM) jointly estimates an agent’s pose and the surrounding 3D scene from sensor observations. In the monocular dense setting, a single RGB camera is used to reconstruct a continuous 3D map, enabling both geometric accuracy and visual realism. The camera poses and dense reconstruction outputs underpin downstream tasks [sengupta2013urban, armeni2019scenegraph, miao2024scenegraphloc, miao2024volumetric, halacheva2024holistic] such as semantic perception, object interaction, and scene editing, and are crucial for applications in VR/AR, robotics, and autonomous driving, where accurate, low-latency 3D perception is essential.
@@ 1.2 Related Work
Classical Visual SLAM    Classical visual SLAM methods can be broadly categorized into two types. The first, similar to incremental SfM [agarwal2011building, schonberger2016structure, liu2024robust], is feature-based SLAM [davison2007monoslam, mur2015orb, mur2017orb, campos2021orb], relying on keypoint extraction and descriptor matching to provide constraints for triangulation and PnP [lepetit2009ep, kneip2011novel] pose estimation. The second, known as direct methods, such as LSD-SLAM [engel2014lsd] and DSO [engel2017direct], optimizes camera poses by minimizing photometric error from pixel intensities while estimating a per-frame depth map. Both categories typically adopt a frontend (feature-based or direct) and a backend for optimization, most often bundle adjustment [triggs1999bundle] to jointly refine poses and structure. However, they rely heavily on accurate camera calibration and are generally limited to sparse 3D reconstructions.
@@
Dense Visual SLAM    To enable denser reconstructions, recent works have incorporated deep learning into either the frontend or the scene representation. For example, DROID-SLAM [teed2021droid] uses RAFT [teed2020raft] for dense optical flow in the frontend and performs dense bundle adjustment on the GPU, while methods such as SuperPrimitive [mazur2024superprimitive] and COMO [dexheimer2024compact] leverage monocular priors (e.g., surface normals, depth distributions) from pretrained models with direct photometric optimization. BA-Track [chen2025back] augments point-based tracking with a scale-grid deformation of monocular depth priors. Neural scene representations have also been adopted, with NICER-SLAM [zhu2024nicer] and MonoGS [matsuki2024gaussian] optimizing camera poses and 3D structure using NeRF [mildenhall2021nerf] and 3D Gaussian Splatting [kerbl20233dgs], respectively. Tracking robustness and geometric accuracy have been further improved by integrating depth priors and auxiliary trackers [zhang2024glorie, sandstrom2025splat, zhang2023go, zhang2023hi, zhang2024hi]. However, most approaches still require accurate camera intrinsics and many struggle to achieve true real-time performance due to the computational demands of dense optimization and neural rendering.
@@
SLAM with 3D Foundation Models    All aforementioned methods require known and accurate camera intrinsics. With the advent of 3D foundation models [dust3r, mast3r, wang2025vggt], several intrinsic-free SLAM frameworks have emerged, aiming to produce dense outputs without calibration. Methods such as Spann3R [wang2025spann3r] and others [liu2025slam3r, wang2025cut3r, cabon2025must3r] extend the two-view DUSt3R [dust3r] model to sequential inputs, directly regressing point clouds in a unified global coordinate system. Reloc3r [dong2025reloc3r] instead regresses only relative poses and performs offline global optimization using SfM techniques [pan2024glomap, sweeney2015theia, zhang2023revisiting]. MASt3R-SLAM [murai2025mast3rslam] extracts dense correspondences from MASt3R [mast3r] and feeds them into a classical optimization pipeline, while submap-based methods [maggio2025vggtslam, deng2025vggtlong] employs the multiview model VGGT [wang2025vggt] to regress local submaps before stitching them via pose graph optimization. While these approaches address some classical limitations, they still face notable drawbacks:
@@
    Current two-view models [dust3r, mast3r] use asymmetric architectures that regress pointmaps of both views to the first view’s coordinates, making it difficult to decouple views for backend optimization (e.g. loop closure).
@@
    Pure regression methods [wang2025spann3r, wang2025cut3r, liu2025slam3r] predict incoming frames with previous memory, but suffer from drift and start forgetting once the trajectory gets longer.
@@
    Methods [wang2025spann3r, wang2025cut3r, liu2025slam3r, murai2025mast3rslam] built on current two-view models inherit the asymmetric architecture with two separate decoders, resulting in large model size. Submap-based methods [maggio2025vggtslam, deng2025vggtlong] employ an even larger multiview model [wang2025vggt] to build submaps, which further increases the size of the frontend model.
@@ 1.3 Contributions of ViSTA-SLAM
To address these concerns, we propose ViSTA-SLAM, a novel real-time monocular visual SLAM pipeline based on symmetric two-view association. At its core is a lightweight Symmetric Two-view Association (STA) model frontend, which takes two RGB images as input and simultaneously regresses two pointmaps in their respective local coordinate frames, along with the relative camera pose between them. During training, we enforce cycle consistency on relative poses and geometric consistency on pointmaps to improve accuracy and stability. Unlike prior 3D models [dust3r, mast3r, wang2025vggt], STA is fully symmetric with respect to its inputs: neither view is designated as a reference, and the same encoder–decoder architecture is applied to both. In the backend, we perform $\mathrm{Sim}(3)$ pose graph optimization with loop closures to mitigate drift and ensure global consistency. To further enhance robustness, each view is represented by multiple nodes rather than a single one, which are connected by scale-only edges to handle scale inconsistencies across different forward passes.
This symmetric design makes our frontend substantially more lightweight than existing methods, with STA being only 64% the size of MASt3R [mast3r] and 35% the size of VGGT [wang2025vggt]. Unlike prior approaches [maggio2025vggtslam, deng2025vggtlong] that group multiple views into a single submap node, our method assigns each view its own nodes in the pose graph. Leveraging the local pointmaps produced by the STA frontend, each node can be represented independently, connected to others solely through relative transformations. Compared to submap-based methods [maggio2025vggtslam], this design yields a more flexible graph structure and greater robustness. The combination of this flexibility and lightweight architecture underpins our choice of a symmetric two-view model as the frontend.
@@
    Our method achieves state-of-the-art performance on the real-world 7-Scenes [shotton2013-7scenes] and TUM-RGBD [sturm12tumrgbd] datasets for both camera trajectory estimation and dense 3D reconstruction.
@@ 2.1 Symmetric Two-view Association Model
Encoder    Our STA model takes in two images $\bm{I}_{i}$, $\bm{I}_{j}$ as input. It uses a shared ViT encoder [dosovitskiy2020image] to divide input images into patches and encode them into features
@@
Symmetric Formulation    Prior approaches [dust3r, mast3r] regress both point maps into the coordinate frame of the first view, thus requiring two separate decoders. In contrast, our model predicts only local point maps and relative poses between views. This fully symmetric formulation makes it possible to use only one decoder. As a result, the number of parameters for decoding is effectively reduced by half (shown in Fig. 4), forming a more compact model for real-time applications. Moreover, producing local view outputs in their own coordinate systems is better suited for the subsequent pose graph optimization, see Sec. 2.3 for details.
Local Point Maps    Given decoding features $\bm{D}_{i/j}^{(b)}$, we use a DPT head [ranftl2021vision] to regress the local point maps $\bm{P}$ and corresponding confidence maps $\bm{W}$:
@@
Since $\bm{M}_{ij}$ is not guaranteed to lie on the $SO(3)$ manifold, we apply SVD orthogonalization [levinson2020analysis] to it to obtain a valid rotation matrix $\bm{R}_{ij}$. Then, the relative transformation is $\bm{T}_{ij}=\left[\bm{R}_{ij}|\bm{t}_{ij}\right]$. With our symmetric formulation, we could also input $\bm{p}_{j}$ into the pose head to regress $\bm{T}_{ji}$. But in practice we only regress one of them for building pose graph.
@@ 2.2 Training Objective
Local Pointmap Loss    Following DUSt3R [dust3r], we apply the confidence-weighted regression loss for all predicted point maps with valid ground truths. Since the reconstruction is up-to-scale, we also normalize the regressed pointmap and the ground-truth pointmap according to their mean Euclidean distance to the origin, $n$ and $\hat{n}$:
@@
Figure 3: Asymmetric vs. Symmetric Architectures. Asymmetric architectures [dust3r, mast3r] use two decoders to regress point maps in a shared coordinate space. our symmetric formulation regresses relative pose and local point maps with only a single decoder, reducing over 36% of the parameters ($\sim$ 0.4 vs. 0.7 billion), while achieving higher accuracy and enabling pose graph optimization in the backend.
@@
NICER-SLAM [zhu2024nicer] 0.033 0.069 0.042 0.108 0.200 0.039 0.108 0.086
DROID-SLAM [teed2021droid] 0.036 0.027 0.025 0.066 0.127 0.040 0.026 0.049
Calib. GlORIE-SLAM [zhang2024glorie] 0.036 0.029 0.014 0.094 0.144 0.053 0.020 0.056
CUT3R [wang2025cut3r] 0.743 0.226 0.363 0.664 0.546 0.381 0.413 0.477
SLAM3R [liu2025slam3r] 0.089 0.048 0.036 0.088 0.196 0.102 0.126 0.098
MASt3R-SLAM [murai2025mast3rslam] 0.063 0.046 0.029 0.103 0.112 0.074 0.032 0.066
VGGT-SLAM [maggio2025vggtslam] 0.037 0.026 0.022 0.103 0.147 0.063 0.095 0.070
@@
Table 1: Camera Trajectory Estimation (ATE RMSE) on 7-Scenes [shotton2013-7scenes]. ViSTA-SLAM performs the best on average.
@@ 2.3 Backend Pose Graph Optimization
Loop Closure    We use Bag of Words [GalvezTRO12] to detect loop candidates, which form new pairs. Then we can feed each candidates pair into our STA model to confirm these proximity. If the predicted confidence score of the relative pose is higher than a predefined threshold $\tau_{p}$, this pair is accepted as a valid loop, and two new nodes connected by a pose edge are added. Each new node is also connected to the first processed node of its corresponding view via a scale edge.
@@
ORB-SLAM3 [campos2021orb] $\times$ 0.017 0.210 $\times$ 0.034 $\times$ $\times$ $\times$ 0.009 N/A
DPV-SLAM [lipson2024dpvslam] 0.112 0.018 0.029 0.057 0.021 0.330 0.030 0.084 0.010 0.076
DPV-SLAM++ [lipson2024dpvslam] 0.132 0.018 0.029 0.050 0.022 0.096 0.032 0.098 0.010 0.054
DROID-SLAM [teed2021droid] 0.111 0.018 0.042 0.021 0.016 0.049 0.026 0.048 0.012 0.038
Calibrated GlORIE-SLAM [zhang2024glorie] 0.128 0.016 0.028 0.021 0.021 0.042 0.020 0.035 0.010 0.036
CUT3R [wang2025cut3r] 0.174 0.592 0.546 0.662 0.467 0.911 0.051 0.845 0.129 0.486
SLAM3R [liu2025slam3r] 0.211 0.861 0.967 0.790 0.755 1.013 0.063 0.986 0.185 0.648
MASt3R-SLAM [murai2025mast3rslam] 0.070 0.032 0.055 0.056 0.035 0.118 0.041 0.116 0.020 0.060
VGGT-SLAM [maggio2025vggtslam] 0.063 0.031 0.048 0.152 0.023 0.133 0.038 0.039 0.020 0.061
@@
Table 2: Camera Trajectory Estimation (ATE RMSE) on TUM-RGBD [sturm12tumrgbd]. ViSTA-SLAM performs the best on average.
@@ 3 Experiments
Evaluation Datasets and Metrics    Following VGGT-SLAM [maggio2025vggtslam], we evaluate our method on standard monocular SLAM benchmarks for camera tracking accuracy and reconstruction quality. We report root mean square error (RMSE) of absolute trajectory error (ATE, in meters) on real-world 7-Scenes [shotton2013-7scenes] and TUM-RGBD [sturm12tumrgbd] datasets using the evo toolkit [grupp2017evo]. Reconstruction quality on 7-Scenes is assessed via RMSE of accuracy, completion, and Chamfer distance (meters), leveraging its ground-truth 3D scenes.
Implementation Details    The frontend STA model is initialized from the weights of DUSt3R [dust3r], and trained on ScanNet [dai2017scannet], ScanNet++[yeshwanthliu2023scannetpp], ARKitScenes[dehghan2021arkitscenes], CO3D [reizenstein21co3d], Aria Synthetic Environments [avetisyan2024ase], and Replica [replica19arxiv] for 7 days using 8 NVIDIA H100 GPUs. We use AdamW optimizer to train our STA model with learning rate $1.5e^{-5}$, weight decay $0.01$, $B=12$, $\alpha^{\text{point}}=0.2$, $\alpha^{\text{pose}}=0.05$, $\lambda_{\text{pmap}}=1$, $\lambda_{\text{pose}}=1$, $\lambda_{\text{gc}}=1$, and $\tau_{p}=0.75$. We conducted evaluations on a machine with an NVIDIA RTX 4090 GPU and an Intel i9-14900KF CPU, with $N=2$ for 7-Scenes and $N=3$ for TUM-RGBD.
Baselines    ViSTA-SLAM is primarily compared with state-of-the-art (SOTA) learning-based SLAM methods in uncalibrated scenarios: VGGT-SLAM [maggio2025vggtslam], MASt3R-SLAM [murai2025mast3rslam], SLAM3R [liu2025slam3r], and CUT3R [wang2025cut3r]. To reduce randomness from VGGT-SLAM’s RANSAC, we run it 5 times per scene as suggested in their paper; results for other methods, including ours, are deterministic. We also compare with SOTA methods using known camera intrinsics [campos2021orb, zhu2024nicer, zhang2024glorie, teed2021droid, lipson2024dpvslam]. Some results are taken from [murai2025mast3rslam, maggio2025vggtslam]. For MASt3R-SLAM and VGGT-SLAM, we keep their original keyframe selection; for ours, CUT3R, and SLAM3R, we use frame strides of 5 (7-Scenes) and 3 (TUM-RGBD). Calibrated methods are shown in gray. Best results are highlighted as first, second, and third.